OpenStudy (anonymous):

Apply Euler's method with the step h = 1/2 to find an approximation of y(1) where y is the solution of the initial value problem y'=1+y y(0)=0

OpenStudy (anonymous):

i get the answer y approximately equals 7/2...can anyone verify?

OpenStudy (anonymous):

\[y(1)=e^1-1=1.71828 \]

OpenStudy (anonymous):

I haven't done Euler's method in a long time. But robtobey's answer doesn't look right to me! I will write down what I have.

OpenStudy (anonymous):

Let \(f(y,t)=1+y\), and we have \(t_0=0\) and \(y_0=0\). Since \(h=\frac{1}{2}\), we need to do two iterations in order to find \(y(1)\). First iteration (finding \(y_1\) which is an approximation of y(0.5): \(f_0=f(0,0)=1\implies y_1=y_0+hf_0=\frac{1}{2}\). Second iteration (Finding \(y_2\), an approximation of \(y(1)\)): \(f_1=f(\frac{1}{2},\frac{1}{2})=1+\frac{1}{2}=\frac{3}{2} \implies y_2=y_1+hf_1=\frac{1}{2}+\frac{1}{2}\times\frac{3}{2}=\frac{5}{4}\). So, \(y(1)\approx \frac{5}{4}\).

OpenStudy (anonymous):

Do you have the final answer?

OpenStudy (anonymous):

The actual solution to the equation is attached and is evaluated at y(1).

OpenStudy (anonymous):

Yeah, it's very easy to find the exact solution from the auxiliary equation \(m-1=0 \implies m=1\), which gives the solution associated with the homogeneous equation to be \(y_c=ce^x\). Its particular solution is some constant, call it \(y_p=A\). Plugging it in the equation yields \(0=1+A \implies A=-1\). So, the general solution is \(y=y_c+y_p=ce^x-1\). And by using the given initial value, we have \(0=c-1 \implies c=1\). So, \(y=e^x-1\). Therefore, \(y(1)=e-1\).